13x+7(-3x-1)=-63 what does x equal?
Found 2 solutions by MathLover1, oberobic:
Answer by MathLover1(20849)
Answer by oberobic(2304)

13x +7(-3x-1) = -63
13x -21x -7 = -63
-8x = -56
x = 7
.
Check the answer to be sure it is correct.
.
13(7) + 7(-(3*7) -1) = ?
91 + 7(-22) = ?
91 - 154 = -63
Correct.
.
Answer: x=7
.
Done.